A Bibliography of Miscellaneous Mathematics Papers

Karl Berry 135 Center Hill Rd. Plymouth, MA 02360

USA

E-mail: karl@cs.umb.edu (Internet)

21 January 2020 Version 1.42

Title word cross-reference

a [LO48]. C^{3} [CJ81, Wic83]. C^{3}CM [Bla82, DS83, Fru81]. C([0, 1]; S′) [Mit84]. D([0, 1]; S′) [Mit84]. L^{2} [BW93, Mal89b]. L^{2}(ℜ) [Aus92]. L^{α} [Hou92]. l_{1} [New65]. M [Vai87]. N [CS92a, BBE89]. N× [CS92a]. ±1 [Saf87]. s [Jac83]. ∑ ±u_{n} [Kah64].

-Channel [Vai87]. -dimensional [BBE89]. -importance [Jac83]. -values [LO48].

1D [LT90].

9th [SIA90].

absojutnaja [Ole81]. absolument [HKK81, KK81]. absolute [KS88a, Szi27]. Achieving [Byr94]. Acoustic [BT93a, Gre88, TH74, Wic92, YWS92, BJ82]. Acoustics [LM83]. activity [Sha85b]. adapted [DW91, DW92]. Adaptive [AS70, Gro76a, Gro76b, LJN^{+}91]. Admiral [Jac]. Advances [Des92]. affine [CS92a, CS92b]. affine-frame [CS92a]. Aids [CJ81]. aléatoirement [Bil65, Bil63]. aléatoires [Kah90, MP76, Mur78]. algebra [Pis79]. algèbrique [Saf87]. algorithm [Pap85, TC83]. Algorithms [Vet84]. Almost [Pey91, Gap77]. along [PK75]. Ambiguities [Fra74, Pat44]. Analog [CG87a, LAG92]. analyse [CM91]. Analysis [BL93, BT93a, CMW92, CW85, Fra74, GS88, HO85, Hou90, MY72, MM79, MZ74, Pat44, RT71, SE90, SW71, Tow81, B^{+}88, Wic91, Wie30]. Analysis/Synthesis [SE90]. Analytic [Bel59, Zah86a, Zah86b]. analytiques [Kah87]. Angle [BLG83, SKL85, TC83]. Antoni [Zyg89]. any [CS92a]. Apparatus [BT93a]. Appl [SIA90]. Application [EG77]. Applications [BL93, BF93, Chu90b, Chu90a, Chu92c, GS58, Pap62, RBC^{+}92, Zah86a, Gap77, Jer77, Kal90]. Approach [Bla82, Hag90, LJN^{+}91, Byr94]. Approximate [Sch62, Wic91]. Approximation [BS92, Cyb89a, Cyb89b, Cyb88, Fre62, HJ90, Mal89b, DPS92, LT90]. Arbitrary [Chu85, Vai87]. Architecture [CG87b]. Arising [Sch63]. Arithmetic [Pis83]. Arrays [Sch60, Byr94]. ART [LJN^{+}91]. ART2 [CG87a, ZdA91]. Aspects [Fla91]. Assessing [DS83]. Assessment [Fow83]. Associative [Spe88]. Asymmetry [Hel72]. Atomes [Per14]. Audio [Vel92]. Auditory [BBN^{+}92, Ben90, BBS92, BT92b, BT93b, Gre88, LAG92, War82, YWS92, BT92a, Sha85a, Sha85b]. Availability [Fow83].

Banach [Gar77, HJ77, HJ74, LT91, MP76, Pis79, Pis89]. band [EG77]. Bandwidth [Chu85]. Banks [Vai92, Vet87]. Based [Hum86, LJN^{+}91, WC92, Yse90]. Bases [Aus92, Mal89b, AWW92, DW92, Dau88, Per91]. basis [DW91, WC92]. bat [SSD92]. be [Kah69]. Bessel [CS92b]. best [WC92]. Between [Hel72]. Beyond [Sch92]. Binary [NK73]. biorthogonal [DW91]. Bit [Vel92, SA85]. Blackwell [RN53]. bodies [Pis89]. Borell [Pis78b]. bornées [Bil65, Bil63, Kah57, KK83]. bounded [CQ92]. Bounds [Byr94]. box [CSW92]. box-spline [CSW92]. Broad [Sch92]. Brownian [DP85, Dav83, Dvo63, OT74, Per83, Tay74]. brownien [Kah76, Kah74, Lév48, HKK81]. browniennes [Kah66]. Burgers [LT90]. Byrnes [Kör80].

C [Pis78b]. Calculating [BA85]. Capability [BLG83]. cardinal [Hig85]. Carlemana [Ole78]. case [DP85]. Category [CG87a]. CELP [SA85]. central [Kot82, Sha85b]. cercle [KK81, Kah82]. Certain [Lit61]. Channel [CEG76, Mal89a, Vai87]. Channels [Chu85]. Chaos [Sch91, CKS91, Kah89]. characterizations [Pis83]. chemical [Kot82]. Circuit [Chu92a]. circular [Kah69]. Clarification [Fow83]. Class [ZdA91, DS52]. Classification [Spe88, YC74, ZdA91]. classifications [Gro76a, Gro76b]. Classify [GS88]. Classifying [LJN^{+}91]. Cochlear [All85, BT93a, PK75]. Code [BA83, SA85, TCW88]. Coder [TCW88]. Codes [CG87a]. Coding [BT93a, EG77, SAH79, SE90, Vel92, VU92, Vet84, WO86, AS79, AS70, AS67, Gro76a]. Coefficients [Byr88, Byr77, LKK77, Kah80, Kwa76, RN53, Saf87]. Cognition [RM88]. coherent [Jac83]. Coifman [AWW92]. Command [Bla82, Dye88]. Communication [Gab46, SW49]. Communications [Bla82]. Compact [BA83]. Compactly [CSW92, CW92, Dau88]. Companies [Off91]. Comparative [ZdA91]. Comparison [BJ82]. Competitive [GG87]. Complete [Byr94]. complex [PW34a, PW34b]. Compression [BT93a, BT92b, BT93b, Chu92a, Wic92]. Computation [DGM90]. Computational [Hag90, AL91]. Computations [LM83]. Computer [DS83, TH74]. Computerized [CJ81]. Computing [Lip87]. Concerns [Jac]. condition [Kah69]. conditions [MP78]. Conference [SIA90]. conjecture [RN53]. Construction [Chu90a, AWW92]. continuability [Zyg33]. continues [Bil65, Bil63, LKK77]. continuity [KS88a]. Continuous [HW89b, HW89a, Rid75]. contribution [Kol24]. Control [Bla82, Dye88, Jur88]. convergence [H89, Gap77, KW87, Kol24, MP78, Pey91, Zyg30]. convergentes [Bil65, Bil63, HKK81, KK81]. convex [Pis89]. Corti [Zwe91]. cosine [AWW92]. Cotype [Pri87, BM85]. Countermeasures [Bla82, CJ81, Wic83]. couples [Saf87]. Coverings [NF77, Kah89]. Crest [Byr94]. Crisis [Dem86]. Criteria [AS79]. criterion [Lev47]. critical [DP85]. Crossings [Hur89, Hum86]. Crystal [Fra74, Pat44]. Curve [RT71]. Curves [DGM90, MM79]. Cuts [Sch92].

D [RN53]. dafür [Ste29]. d’après [Pis78b]. daßder [Ste29]. Data [BT93a, BT92b, BT93b, BM83, CV86, Chu92a, Chu90b, BJ82]. Decentralized [Tsi86]. Decimated [Vai87]. decimation [CEG76]. Decision [CJ81, RN87, TVB87]. Decisions [Zah87]. Decoding [BT93a]. decomposition [CEG76, Pap85]. Decompositions [AH92, Mal89a]. Deming [GG87]. Density [Lev40]. Dependency [BA85]. Design [Chu85, Fru81, Vai87, Lev47]. Designing [Bla82]. Detection [CV86, MH80, RT71, Tsi86, Van68, MH92]. Detectors [MT72, Gro76a]. development [Gro76a]. Deviations [BHxx]. Dichotomie [BM85, Pri87]. dictionaries [MZ92]. Differential [Wie23, DW92, HKRX91, KW84, Kal90, KS88b, Lia90, LM91, Per91, Wal86]. Diffusion [KMW90]. diffusions [KS88b, Kot82]. Digital [Coh92, GW87, SAH79, VD89, VU92]. Dilation [Aus92]. dimension [GLRT90, Per83, Sal50]. dimensional [BBE89, KW84, Kal90, SKL85, Wei 2]. Dimensions [SA90]. Directed [Hag90]. disc [Mur81, Off72]. Discrete [Chu90b, HW89b, HW89a, MT72]. Distributed [RN87, RM88, KW84]. Distribution [Kah87, PK75, LO48, Mur78, Mur81, Off72]. distributions [Mit81]. does [OT74]. domain [PW34a, PW34b]. driven [HKRX91, KPA88]. droite [Mal59]. duality [CW92]. duals [CKS91, KMW90]. dynamics [Wei 2].

Ear [SAH79, Sch74, SSD92]. Edge [MH80, RT71]. Effectiveness [BA85, DS83]. Efforts [Off91]. Eigenschaft [Ban30]. element [Yse90]. élémentaires [IIK65]. elements [Jac83]. elliptic [GLRT90, Jaf90, Tch91]. Emphasis [Fla91]. Empirical [BM83]. Employing [Byr88]. Engr [SIA90]. Ensembles [Pis78a, Dru70, Kah66]. Entropy [WC92, Ben87, Pap81, Wei 2]. Equal [Byr88, Chu85]. Equation [TH74, CKS91, KW84, LT90, SKL85, TC83]. equations [HKRX91, KPA88, Kal90, KMW90, KS88b, KR81, Lia90, LM91, Per91, Wal86]. ergodic [Hou92]. Error [AS79, Beu66, RHW88, Lev47]. Espaces [Pri87, BM85, MP76]. Estimates [BHxx]. Estimation [Ben91, Van68, YC74, Zah86b, Kav79, Pap81, Pap85]. étude [Kol24]. Euclidean [SW71]. Evaluation [Bla82, ZdA91]. everywhere [Gap77, Pey91]. evoked [Sha85b]. evolution [KPA88, KR81]. Excited [TCW88, SA85]. expansions [DGM86, YT67]. expectation [Gro76b]. Experimental [MM79, BJ82]. Exploiting [SAH79]. Explorations [RM88]. Extensions [SA90, Jer77]. Extremal [New65, Sha51, Saf87].

Face [IRY92]. Factor

[Aus92, Byr94, Wic91]. fail [OT74]. Fast [Wic91, Jaf90]. feature [Gro76a]. Feedback [Gro76b]. fenêtre [CM91]. fiber [PK75]. Field [LM83]. Fields [ZH84]. Filter [Byr88, Chu85, Vai92, Vet87, Lev47]. Filters [BT93a, EG77, SA90, Vai87, Byr94]. Finding [Zwe91]. finite [Yse90]. Five [Hig85]. fonctions [LKK77, Kah57, Kah87, KK83]. Forms [GS58]. Formula [BBE89, Ben91]. Formulas [Chu90a]. Foundations [AR88]. Fourier [CM91, You80, Bil65, Bil63, LKK77, DS52, Fre66, Hou90, HKK81, KK81, IIK65, Kah86, KK83, Kah82, Kol24, Ole81, PW34a, PW34b, Pap62, Saa79, SW71]. Fourierreihen [Szi27]. Fouriersche [Fre62]. Fractals [Sch91]. frame [CS92a]. Frames [BW93, BL93, Ben92a, CS92b]. framework [CD91]. free [Beu66]. Frequency [CGT89, Fla91, MZ92]. Freud [Kah86]. Function [Cyb89a, Cyb89b, Cyb88, FJW91, LO48, Off72]. Functional [M^{+}88, Ost93]. Functions [BS70, Byr94, Gap77, Kah85, PZ32, PWZ33, PZ30, Rid75, Sal50]. funktsii [Ole78, Saa79]. Fusion [CV86, Hag90, Per87, TVB87].

Gabor [BW93]. Galerkin [DW91, Gop 2]. Gap [Lev40]. gaps [Fre66, Rud60]. Gaussian [Kah83, MT72]. Gaussiennes [Kah87]. Gaussiens [Pis78a]. Gaussovskie [Hru81]. general [CD91]. Generalized [Benxx, Wie30]. genetics [FV81]. géométrie [MP76]. geometry [Pis89]. Geza [Kah86]. GHz [AL91]. given [Sal50]. gladkosti [Saa79]. Grenze [Ste29]. Growing [NK73]. Guide [GG87, Inc89].

Hadamard [Fre66]. Handbook [Jur88, Sko70]. harmonic [Wie30]. Hausdorff [Per83, Sal50]. Having [Vai87]. Hearing [LMNY82, Pic88, Sch75]. Hidden [GS88]. Hierarchy [Zah86a, Zah86b]. High [Kav79, SA85]. Homéomorphismes [KK81, Kah82]. homogeneous [Pis79]. Homometric [RS82]. Human [MZ74, SAH79]. Hybrid [IRY92]. hydrodynamics [Wei 2]. hyperbolic [GLRT90].

IFD [BLG83]. ihre [Ste29]. II [Gro76b, Inc89, Sha85b, ZdA91, Zyg68]. illusions [Gro76b]. Image [BA83, GW87, KJ55, MZ74]. Images [Kah66, LJN^{+}91, Mal89a, WO86, ZdA91]. Impedance [Zwe91]. importance [Jac83]. impossibilité [Mal59]. Improve [Off91]. indépendantes [MP76]. independent [HJ74]. inégalités [Pis78b]. Inference [Gib71]. Infinite [KW84, Kal90, Sch91]. Information [ZH84]. Infra [ZdA91]. Infra-red [ZdA91]. Infrared [LJN^{+}91]. Inhibiting [MY72]. inhibition [Sha85b]. Input [CG87a]. Instant [BS70]. instants [Kah76]. Integral [Pap62, LO48]. Integral’nie [Saa79]. integrals [Gap77]. Interface [LM83]. Internal [RHW88].

International [SIA90]. Interpolation [Chu90b, Chu90a, HJ90, CD91, CEG76, Pir67]. interpolation/decimation/tree [CEG76]. interpolatory [YT67]. Interpretation [Sab78]. interval [CQ92]. Introduction [Chu92b, Hov88, Lip87, Pic88, SW71, You80, Wal86]. Invariants [Pri87, BM85]. Irregular [Ben92a, BT93a, BBN^{+}92, Ben90, BBS92]. irrégularité [Kah74]. irregularities [Tay74]. irregularly [Beu66]. ispravlenii [Hru81]. iterated [OT74].

J.S. [Kör80]. Jordan [Hal91, Jac].

Judgment [CW85].

Kahane [Pis78b]. Kbps [TCW88].

Khintchine [Pis78b]. Khintchine-Kahane [Pis78b]. koeffitsienti [Saa79]. Konvergenz [Szi27]. Konvergenzkreis [Ste29].

lacunaires [H89, IIK65]. lacunary

[Kah86, Pey91, Zyg30, Zyg32]. lakunären [Ban30]. Language [LMNY82]. Laplacian [BA83]. Lapped [Mal92]. Large [BHxx, Kot82]. Lateral [Sha85b]. Law [Kot82, OT74]. Laws [Sch91]. Layered [GS88]. Learning [RHW88]. lecons [Kah82]. Lectures [Dau92]. Levinson [Pap85]. limit [Kot82]. line [Ben87]. Linear [Hou90, NB70, Sch60, TCW88, B^{+}88, GLRT90, SA85]. Linguistic [MM79]. Littlewood [FJW91]. Littlewood-Paley [FJW91]. Local [AWW92, CD91]. locale [Kah74]. logarithm [OT74]. Long [TH74]. Long-Range [TH74]. loss [BJ82]. Lossless [Chu92a]. Losslessness [VD89]. Low [Byr94, SA85]. Lücken [Szi27].

Machine [CG87b]. Magnitude [Byr88]. Making [CJ81]. Malliavin [Kah59]. Management [Sch92, Off91]. Manufacturing [Sch92]. Mapping [Spe88]. Markov [FV81]. Martingales [Mit81, KPA88]. Masking [Hel72, SAH79]. Massively [CG87b]. Matching [MZ92]. Mathematical [Sch60, SW49]. Mathematics [BF93]. mathématiques [Sal67]. Matrices [CW85]. Maximally [Vai87]. Maximum [Pap81, Ben87]. McKean [CKS91]. McKean-Vlasov [CKS91]. Mean [Lit61]. measurable [Gap77]. measure [FV81, KS88b]. measure-valued [FV81, KS88b]. measures [HKRX91, KS88a]. Memories [Sch63]. Memory [Spe88]. Menchoff [Kat75]. Men’sova [Hru81]. Message [Ost93]. Meta [HO85]. Meta-Analysis [HO85]. Method [BT93a, Chu92a, TH74, Zah86a, Gop 2, Per91]. Methods [BS92, BM83, CGT89, Fla91, HO85, SIA90, Sch62, Zah86b, DW91, DPS92, Jaf90, WC92]. metric [Kah69]. Meyer [AWW92]. Microstructure [RM88]. Millimeter [AL91]. minimum [Ben92b]. Minutes [Sch91]. Mirror [Chu85, EG77, SA90, Vai87, Byr94]. Model [Ben90, BA85, LM83, LAG92, Wal81]. modeling [All85]. Models [BBN^{+}92, BBS92, BT92b, BT93b, DS83, Mal89a, Sch75, BJ82, KW84, KW87]. Modern [VD89]. Modulation [Van68]. moduli [Saa79]. Modulus [Byr77, Ben92b]. MOEs [Wic83]. Moment [Gop 2]. monotonic [Sal50]. motion [Dvo63]. motivated [BT92a]. Mouvement [HKK81, Kah74, Kah76, Lév48]. moyenne [Kah57]. moyenne-périodiques [Kah57]. Multi [ZdA91]. Multi-Class [ZdA91]. Multidimensional [M^{+}88, Vet84]. Multifrequency [Mal89a]. Multilevel [DK92, BPX90, Yse90]. Multiple [CV86, Per87, SA90, TVB87]. multiplication [Kah89]. multiplicative [Kah89]. multiplicity [Sal44]. Multiprocessors [Ost93]. Multiqueuing [Sch62]. Multirate [Vai92, Vet87]. Multiresolution [BL93, Mal89b, VU92].

Multiscale [AH92]. Multivariate [HJ90]. mutual [KS88a]. Mutually [MY72].

natürliche [Ste29]. NAVSEA [Fow83].

Necessary [MP78]. nerve

[PK75, Sha85a, Sha85b]. Nets [Lip87]. Network [DAR90, GS88, IRY92, LJN^{+}91]. Networks [MY72, Per87, Spe88, ZdA91]. Neural [CG87b, DAR90, IRY92, Lip87, LJN^{+}91, Per87, Spe88, ZdA91, Gro76a, Wal81]. NeuralWorks [Inc89]. Neurocomputing [AR88, HN90]. Neurodynamics [Ros59]. neuronal [KW87]. Neurons [MY72, KW84]. neustranimymi [Ole78]. Noise [BT93a, Hel72, MT72]. noisy [Kav79]. non [Jac83, MT72]. non-coherent [Jac83]. non-Gaussian [MT72]. Nonharmonic [You80, DS52]. Nonlinear [BT93a, BS, GLRT90]. nonorthogonal [DGM86]. Nonparametric [Gib71]. Nonseparable [SA90]. Nonstationary [Fla91]. nonuniform [YT67]. Notch [Byr88]. Note [NK73]. Notes [PWZ33]. Nuclear [HKRX91, CKS91, KPA88, KMW90]. Num [SIA90]. Number [Chu85]. numbers [Kot82]. Numerical [BS92, Lia90].

ob [Hru81]. oblate [AL91]. Ocean [LM83, TH74]. odd [CS92a]. Oeuvres [Bor72, Sal67]. Offensive [Hal91, Wic83]. often [OT74]. olfaction [Gro76b]. Ondelettes [Mey90]. One [Byr77, GLRT90]. Opérateurs [Mey90]. Operational [Fow83, Fru81]. operators [DW92, DPS92, Tch91]. Optimal [CV86, RN87, TVB87, Byr94]. Optimizing [SAH79]. ordinary [LM91]. Organ [Zwe91]. Organisation [CG87a]. Organising [CG87b]. Orthogonal [Mal92]. Orthonormal [Dau88, Mal89b, Byr94]. oscillation [Dvo63]. osobennostjami [Ole78]. oversampling [CS92a].

Packets [BS, Wic92]. Painless [DGM86]. Paley [FJW91]. papers [Zyg89]. Parabolic [TH74, GLRT90, SKL85, TC83]. Paradise [Sch91]. Parallel [BPX90, CG87b, Ost93, RM88, Gro76a]. parfaits [Kah66]. partial [KS88b, Lia90, Per91, Wal86]. partition [PK75]. partout [H89]. Parts [ZdA91]. Passing [Ost93]. path [OT74, Tay74]. Pathology [Sch74]. Pattern [CG87b, YC74, Gro76a, Gro76b]. Patterns [CG87a]. Paul [Kah59]. Perception [MZ74, War82]. peremennoy [Ole81]. Perfect [CEG76, Vai87]. Perfect-reconstruction [Vai87]. Performance [RN87, Off91]. périodiques [Kah57]. Periphery [Gre88, LAG92]. Phase [CGT89, Hur89]. Physiology [Pic88]. pjada [Ole81]. Plancherel [BBE89, Ben91]. Planning [Hag90]. plicants [Jac83]. PNN [ZdA91]. points [DP85, Dav83, Kah83, Per83]. pointwise [Hou92]. Poisson [HKRX91]. Policy [Fow83]. polynômes [Kah80, Saf87]. polynomial [Kör80]. Polynomials [BHxx, Byr77, Lit61, New65, Sha51, Ben92b]. population [FV81]. Position [GG87]. potentiel [Kah90]. Potenzreihe [Ste29]. Pour [Pri87, BM85]. Power [Sch91, Sha51, RN53, Zyg33]. Practice [Off91]. preconditioners [BPX90, Yse90]. preconditioning [DK92]. prediction [B^{+}88, Lev47, SA85]. Predictive [AS79, AS67, TCW88, AS70]. preserves [CS92a]. presque [H89]. prime [Jac83]. Principle [Byr94, CW92]. Principles [Ros59, Sab78, Uri83]. Probabilistic [Spe88]. probabilities [Mit84]. Probability [HJ77, LT91, Pap91]. Problem [New65, Sch62, Sch63, Pir67]. Problème [Saf87]. Problems [Erd57, NB70, Sha51, GLRT90, Jaf90]. Proc [SIA90]. Process [Zah86a, Zah86b, Dvo63]. Processes [Pap91, Bel59, FV81, Kah83, Pir67]. Processing [Coh92, CMW92, Fla91, GW87, KJ55, Mal92, RV91, RM88, VD89, MH92, Sha85a, Sha85b]. Processus [Lév48, Pis78a]. product [KS88a]. products [Bou85, Kah91, Pey91]. produits [H89]. Professional [Inc89]. Programming [Ost93, Zah87]. Propagation [AL91, CKS91, RHW88, TH74]. Properties [SAH79, SZ54, YT67]. property [Vai87]. protsessi [Hru81]. pseudo [DW92]. pseudo-differential [DW92]. pseudodifferential [DPS92]. Pursuits [MZ92]. Pyramid [BA83].

Quadrature

[Chu85, EG77, SA90, Vai87, Byr94]. Qualitative [Zah87, Ben87]. Quality [GG87, Jur88, Sch92, Off91, SA85]. Quatre [Kah82]. Question [Hal91]. Queueing [Sch63].

Radar [Sko70]. Rademacher [Kwa76]. RADM [Hal91]. Rain [AL91]. raindrops [AL91]. ralentissement [Kah76]. Random [BHxx, Kah89, NF77, Pap91, Bel59, Gar77, HKRX91, HJ74, Kah91, Kah85, LO48, MP78, Mit81, Mur81, Off72, PWZ33, Pir67, RN53, SZ54]. Randomly [Rid75]. Range [TH74]. rapport [H89]. Rates [Vel92, SA85]. Rational [Aus92]. ray [Fra74, Pat44]. reactions [Kot82]. real [Ben87]. Rear [Jac]. Receptive [ZH84]. Recirculating [Sch63]. Recognition [CG87a, CG87b, IRY92, YC74]. Reconstruction [BT93a, Vai87]. recording [Gro76a, Gro76b]. Recouvrements [Kah90]. recovery [Beu66]. red [ZdA91]. Regularity [Tay74]. regularized [LT90]. Reihen [Ban30, Fre62]. Related [BW93]. remarkable [Pis79]. remarque [Mur78]. Remarques [CM91]. Remote [Sab78, Tow81]. Representation [LAG92, ZH84, BT92a, Sha85a]. Representations [Hum86, RHW88, YWS92]. Research [AR88]. Resolution [LT90, Jaf90, Kav79]. Resonance [LJN^{+}91]. respect [Pey91]. response [Wal81]. responses [PK75, Sha85a]. Retrieval [Hur89]. Review [VU92, Jer77, Pap81]. Riesz [Bou85, H89, Kah91, KS88a, Pey91]. RMS [Lev47]. Role [VD89]. Rudin [Benxx, Saf87]. Rudin-Shapiro [Benxx, Saf87]. Rules [Tsi86]. Ryadi [Bar61].

s [Ole78]. Salem [Saf87]. samples [Beu66]. Sampling [Ben92a, BT93a, BBN^{+}92, Ben90, BBS92, B^{+}88, Jer77, YT67]. Satzes [Szi27]. Scale [Fla91, Hum86, BT92a]. Scaling [M^{+}88]. Scattering [AL91]. Scene [RT71]. Schemes [EG77]. Sciences [SIA90]. Search [NB70]. Selected [Zyg89]. Selection [Zah87, WC92]. Self [CG87a, CG87b]. Self-Organisation [CG87a]. Self-Organising [CG87b]. Sensing [Sab78, Tow81]. Sensor [CV86, Hag90, Hov88, Per87, RN87, TVB87]. Sequences [Byr94, CS92b, Gap77]. Series [Sha51, You80, Zyg68, DS52, Fre66, Hig85, Kah86, Kah85, Kwa76, MP78, Mur81, PZ32, PZ30, Pey91, Rud60, RN53, SZ54, Zyg33, Zyg30, Zyg32, Bor97, Bor96, H89, HKK81, KK81, IIK65, Kah82, Kol24, Mur78, Bil65, Bil63, KK83, MP76]. set [Kah69]. Sets [RS82, Sal44, Bou85, Kah91, Pis83, Rid75]. SEW [Hal91]. Shannon [Jer77]. Shapiro [Benxx, Saf87]. Ship [ZdA91]. shodimost [Ole81]. short [Hig85]. Sidon [Bou85, Dru70, Pis83, Pis78a, Rid75]. Sigmoidal [Cyb89a, Cyb89b, Cyb88]. Signal [AH92, Coh92, CMW92, Fla91, Mal92, RV91, VD89, Wic92, BT92a, B^{+}88]. Signals [BT93a, MT72, YWS92, AS79, AS70, AS67, Beu66, Kav79]. signs [SZ54]. Silicon [LAG92]. Simulation [MY72, TH74, Zah86b]. sine [AWW92]. singular [Sal50]. Singularity [MH92, KS88a]. SLLN [Hou92]. Slow [Kah83, DP85, Dav83, Per83]. smooth [AWW92]. solution [GLRT90, Lia90]. solving [Per91]. Some [Erd57, Fla91, FV81, Kah85, SZ54, Vet84, KS88b, PZ32, PZ30, YT67]. sommes [Kah64]. Sonar [GS88]. Sound [Uri83]. sounds [Sha85a]. Source [Vel92]. Space [CGT89, Hum86, Gar77, GLRT90, HKRX91, HJ74, KPA88, Pis89, Wie23]. space-valued [KPA88]. spaced [Beu66]. Spaces [BW93, FJW91, SW71, CKS91, HJ77, KMW90, LT91, Yse90]. spatial [LT90]. spatially [KW84]. Specification [Fow83]. Spectral [Ben91, Hou92, Kav79, Pap81, Pap85]. spectrale [Mal59]. spectrum [Sal50]. Speech [LMNY82, LAG92, SAH79, Sha85a, Sha85b, AS79, AS70, AS67, SA85]. Spherical [Ben91, AL91]. spheroidal [AL91]. Spline [LM91, CW92, CSW92]. Splines [Chu90b]. Split [EG77, TC83]. Split-band [EG77]. split-step [TC83]. Splitting [CEG76, Yse90]. stability [YT67]. Stable [CG87a]. Statistical [Gib71, HO85]. step [TC83]. Stochastic [Hou90, KPA88, KS88b, KR81, Pap91, Gap77, HKRX91, KW84, Kal90, KW87, Wal81, Wal86]. stochastiques [Lév48]. stories [Hig85]. Structural [BM83]. Structure [RS82, Saf87]. Structures [Fra74, Pat44]. Study [DAR90, FJW91, Zah86b, AL91]. Subband [SE90, Vet84, WO86]. Subjective [AS79, CW85]. Sucectvovanie [Ole78]. sufficient [MP78]. Sums [Gar77, HJ74]. Superpositions [Cyb89a, Cyb89b, Cyb88]. superpozitsii [Saa79]. Supervised [IRY92]. Supervised/Unsupervised [IRY92]. supported [CW92, CSW92, Dau88]. Suppression [BT93a]. Surfaces [DGM90]. Survey [Zah86a]. Swath [Sch92]. synthèse [Mal59]. Synthesis [SE90]. System [BA85, Sha85a, Sha85b]. Systems [CV86, Hov88, RN87, TVB87, Benxx, Jac83].

T&E [Bla82]. Targets [GS88, LJN^{+}91]. Task [Hag90]. Taylor [Bor97, Bor96, Mur78, Mur81]. Techniques [SE90, VU92, CEG76]. Technology [Des92]. Television [VU92]. Ten [Dau92]. Teorema [Hru81]. terms [SZ54]. Terrain [Tow81]. Test [Bla82]. Tests [Fru81]. Their [Chu90b, GS58, RBC^{+}92]. theorem [Ben87, B^{+}88, Gap77, Gop 2, Hou92, Jer77, Kat75, Kot82, Kwa76]. théorème [Kah59]. Théorèmes [IIK65]. Theorems [Lev40, NF77]. théorie [Kah90]. Theory [BS92, Chu92c, FJW91, Gab46, LJN^{+}91, MH80, Sch60, SW49, Vai87, Van68, Vet84, FV81]. three [BJ82, SKL85]. three-dimensional [SKL85]. Threshold [Tsi86]. tight [CS92a]. Tightness [Mit84]. Time [CGT89, Fla91, MT72, BT92a, MZ92]. Time-Frequency [CGT89, Fla91, MZ92]. Time-Scale [Fla91, BT92a]. Toeplitz [GS58]. Tone [Hel72]. Tool [BA85]. Total [Sch92]. Trained [GS88]. Transduction [Gre88]. transfer [Wei 2]. Transforms [HW89b, HW89a, Mal92, PW34a, PW34b]. transmission [BJ82]. Treatment [BM83]. tree [CEG76]. Trees [NK73]. Trigonometric [BHxx, Lit61, Rud60, Zyg68, Ben92b, MP78, Pey91, SZ54, Zyg30, Zyg32]. Trigonometriceskie [Bar61]. trigonométriques [H89]. trigonometrische [Fre62]. trigonometrischen [Ban30]. Tutorial [Chu92c, Jer77]. Two [Yse90, Wei 2].

u [Saa79]. U.S [Off91]. Uncertainty [Byr94]. Underwater [Uri83]. Unified [Byr94]. uniform [MP78]. uniformément [Bil65, Bil63]. unimodulaires [Kah80]. uniqueness [Kah69, Sal44]. unit [Mur81, Off72]. Units [GS88]. universal [Gro76a, Gro76b]. Unsolved [Erd57].

Unsupervised [IRY92]. use [CEG76]. Useful [Wic83]. Users [Inc89]. Using [BT93a, IRY92, TH74, LT90, Per91].

valeurs [Kah87, Mur78]. Value [Mur81]. valued [FV81, Gar77, HKRX91, HJ74, KPA88, KS88b]. Values [Lit61, LO48, Off72]. Variables [Pap91, Gar77, HJ74, MP76]. various [Jer77]. vector [Kwa76]. vectorielles [Kah64, MP76]. Verallgemeinerung [Szi27]. Vertex [Chu90b]. very [SA85]. Visual [RT71, ZH84]. Vlasov [CKS91]. Voice [EG77]. Voiced [LAG92]. Voiced-Speech [LAG92]. volume [Pis89].

Wahrscheinlichkeit [Ste29]. Walsh [BS70]. Warfare [Dye88]. Wave [AL91, LM83, SKL85]. Wavelet [Aus92, BT93a, BS, BBN^{+}92, Ben90, BBS92, BT92b, BT93b, CMW92, DW91, DW92, DPS92, GLRT90, HW89b, HW89a, Jaf90, Mal89a, Mal89b, Wic92, Gop 2, LT90, Per91]. Wavelet-Galerkin [DW91, Gop 2]. Wavelets [BF93, Chu92b, CQ92, Chu92c, Coh92, CGT89, Dau92, Lia90, RV91, RBC^{+}92, Vai92, Wei 2, AWW92, CW92, CSW92, Dau88, LM91, MH92, Tch91]. Weak [KW87]. whose [SZ54, Sal50]. Wide [BLG83, SKL85, TC83]. wide-angle [SKL85, TC83]. Wiener [BBE89, Ben91, Lev47]. Wiener-Plancherel [BBE89, Ben91]. Wold [Pap85].

X [Fra74, Pat44]. X-ray [Fra74, Pat44].

Zamena [Ole81]. Zero [Hum86, Hur89]. Zero-crossings [Hum86]. zeros [LO48, Kah76]. Zygmund [Zyg89].

References

Akansu:MSD-92

[AH92] Ali Akansu and R. Haddad. Multiscale Signal Decompositions. Academic Press, New York, 1992.

Aydin:IGRS-29-593

[AL91] K. Aydin and Y. Lure. Millimeter wave scattering and propagation in rain: A computational study at 94 and 140 GHz for oblate spheroidal and spherical raindrops. IEEE Trans. Geo. Remote Sensing, 29(4):593–601, July 1991.

Allen:CM-2-85

[All85] J. Allen. Cochlear modeling.

ASSP Magazine, 2:3–29, 1985.

Anderson:NFR-88

[AR88] J. Anderson and E. Rosenfeld.

Neurocomputing: Foundations of Research. MIT Press, Cambridge, MA, 1988.

Atal:PC-67-360

[AS67] B. Atal and M. Schroeder. Predictive coding of speech signals. Proc. 1967 Conference on Communications and Processing, pages 360– 361, November 1967.

Atal:BST-49-1973

[AS70] B. Atal and M. Schroeder. Adaptive predictive coding of speech signals. Bell Sys. Tech. J, 49:1973– 1986, 1970.

Atal:ASSP-27-247

[AS79] B. Atal and M. Schroeder. Predictive coding of speech signals and subjective error criteria. IEEE Trans. ASSP, 27:247–254, 1979.

Auscher:WTA92-439

[Aus92] P. Auscher. Wavelet bases for L^{2}(ℜ) with rational dilation factor. In Ruskai et al. [RBC^{+}92], pages 439–452.

Auscher:WTTA-92-237

[AWW92] P. Auscher, G. Weiss, and M. V. Wickerhauser. Local sine and cosine bases of Coifman and Meyer and the construction of smooth wavelets. In Chui [Chu92c], pages 237–256.

Butzer:STL-JDM-90-1

[B^{+}88] P. Butzer et al. The sampling theorem and linear prediction in signal analysis. Jber. Deutsch. Math.—Verein, 90:1–70, 1988.

Burt:ICOMM-31-562

[BA83] P. Burt and T. Adelson. The Laplacian pyramid as a compact image code. IEEE Trans. Comm., COM-31:532–540, 1983.

Byrnes:IREL-85-17

[BA85] J. S. Byrnes and Robert C. Angell. The dependency model: A tool for calculating system effectiveness. IEEE Trans. Reliability, R-34(1):17–23, April 1985.

Banach:SM-2-207

[Ban30] S. Banach. Uber eine eigenschaft der lakunären trigonometrischen reihen. Studia Math., 2:207–220,

Bari:TCR-61

[Bar61] Nina Bari. Trigonometričeskie Ryadi. Gos. izd-vo fizikomatematicheskoi lit-ry, Moscou, Russia, 1961.

Benedetto:AAP-10-457

[BBE89] J. Benedetto, G. Benke, and W. Evans. A n-dimensional Wiener-Plancherel formula. Adv. Appl. Math, 10:457–487, 1989.

Benedetto:PRO-92-2

[BBN^{+}92] J. J. Benedetto, J. S. Byrnes, D. J. Newman, H. S. Shapiro, and A. Teolis. Wavelet auditory models and irregular sampling. Technical Report AR-92-2, Prometheus Inc., 1992. Annual Phase II report.

Benedetto:WAM-92-2

[BBS92] J. Benedetto, J. S. Byrnes, and H. S. Shapiro. Wavelet auditory models and irregular sampling. Monthly status report, February 1992.

Belayev:TPA-4-402

[Bel59] Yu.K. Belayev. Analytic random processes. Theor. Probability Appl., 4:402–409, 1959.

Benedetto:IEPT-10-761

[Ben87] J. Benedetto. A qualitative maximum entropy theorem for the real line. Integral Equations and Operator Theory, 10:761–779, 1987.

Benedetto:Prometheus-90-1

[Ben90] J. Benedetto. A wavelet auditory model and irregular sampling. Technical Report 90-1, Prometheus Inc., 1990.

Benedetto:SMA-22-1110

[Ben91] J. Benedetto. The spherical Wiener-Plancherel formula and spectral estimation. SIAM J. Math. Anal., 22:1110–1130, 1991.

Benedetto:ISF-92-445

[Ben92a] John Benedetto. Irregular sampling and frames. In Chui [Chu92b], pages 445–507.

Benke:PAMS-113-757

[Ben92b] G. Benke. On the minimum modulus of trigonometric polynomials. Proc. Amer. Math. Soc., 114(3): 757–761, March 1992.

Benke:JFAA-ta

[Benxx] G. Benke. Generalized Rudin-Shapiro systems. Journal of Fourier Analysis and Applications, 19xx. To appear.

Beutler:SIAM-8-328

[Beu66] F. Beutler. Error free recovery of signals from irregularly spaced samples. SIAM Review, 8:328–335, 1966.

Benedetto:WMA-93

[BF93] J. Benedetto and M. Frazier, editors. Wavelets: Mathematics and Applications. CRC Publishers, Boca Raton, FL, 1993.

Benke:ELD-pre

[BHxx] George Benke and W. J. Hendricks. Estimates for large deviations in random trigonometric polynomials, 19xx. Preprint.

Billard:SM-22-309

[Bil63] P. Billard. Séries de Fourier aléatoirement bornées, continues, uniformément convergentes. Studia Math., 22:309–329, 1963.

Billard:ASENS-82-131

[Bil65] P. Billard. Séries de Fourier aléatoirement bornées, continues, uniformément convergentes. Ann. Scient. Ec. Norm. Sup., 82:131– 179, 1965.

Boyles:JHAPLTD-3-67

[BJ82] C. A. Boyles and G. W. Joice.

Comparison of three acoustic transmission loss models with experimental data. Johns Hopkins APL Tech. Dig., 3:67–76, 1982.

Benedetto:ICASSP-93-ta

[BL93] J. Benedetto and S. Li. Multiresolution analysis frames with applications. In Proceedings of ICASSP, 1993. To appear.

Blanche:MTR-24155wp

[Bla82] Raul C. Blanche. An approach to designing a command, control and communications countermeasures test and evaluation (C^{3}CM T&E). Working Paper 24155, Mitre Corporation, July 1982.

Botseas:NUSC-83-6905

[BLG83] George Botseas, Ding Lee, and Kenneth E. Gilbert. IFD: Wide angle capability. Technical Report 6905, NUSC, 1983.

Braverman:SME-83

[BM83] E. M. Braverman and I. B. Muchnik. Structural Methods for Empirical Data Treatment. Nauka, Moscow, USSR, 1983.

Bourgain:CRASP-300-263

[BM85] J. Bourgain and V. Milman. Dichotomie du cotype pour les espaces invariants. C.R. Acad. Sci. Paris, 300:263–266, 1985.

Borel:CRASP-123-1051

[Bor96] E. Borel. Sur les séries de Taylor. C.R. Acad. Sci. Paris, 123:1051– 1052, 1896.

Borel:AM-20-243

[Bor97] E. Borel. Sur les séries de Taylor. Acta Math, 20:243–247, 1897.

Borel:O-72

[Bor72] E. Borel. Oeuvres. CNRS Paris, 1972. Tome I.

Bourgain:AIF-35-137

[Bou85] J. Bourgain. Sidon sets and Riesz products. Ann. Inst. Fourier (Grenoble), 35(1):137–148, 1985.

Bramble:MC-55-1

[BPX90] J. H. Bramble, J. E. Pasciak, and J. Xu. Parallel multilevel preconditioners. Math. Comp., 55:1–22, 1990.

Benedetto:NWP-ip

[BS] J. Benedetto and S. Saliani. Nonlinear wavelet packets. In preparation.

Byrnes:SR-12-131

[BS70] J. S. Byrnes and D. A. Swick. Instant Walsh functions. SIAM Review, 12:131, 1970.

Braess:NMAT-92

[BS92] D. Braess and L. L. Schumaker, editors. Numerical Methods of Approximation Theory. Birkhäuser Verlag, Basel, 1992. To appear.

Benedetto:STFS-92

[BT92a] J. Benedetto and A. Teolis. An auditory motivated time-scale signal representation. In IEEE-SP International Symposium on Time-Frequency and Time-Scale analysis. Victoria, BC, 1992.

Benedetto:WAMDC-92

[BT92b] J. Benedetto and A. Teolis.

Wavelet auditory models and data compression. Preprint, 1992.

Benedetto:NMA-93

[BT93a] J. Benedetto and A. Teolis. Nonlinear method and apparatus for coding and decoding acoustic signals with data compression and noise suppression using cochlear filters, wavelet analysis, and irregular sampling reconstruction. U.S. Patent pending, 1993.

Benedetto:WAMDC-93

[BT93b] J. Benedetto and A. Teolis.

Wavelet auditory models and data compression. Applied and Computational Harmonic Analysis, 1993. To appear.

Benedetto:GFRS-93

[BW93] J. Benedetto and D. Walnut. Gabor frames for L^{2} and related spaces. In Benedetto and Frazier [BF93].

Byrnes:BLMS-9-171

[Byr77] J. S. Byrnes. On polynomials with coefficients of modulus one. Bull. London Math. Soc., 9:171– 176, 1977.

Byrnes:ASSP-36-1783

[Byr88] J. S. Byrnes. A notch filter employing coefficients of equal magnitude. IEEE Trans. ASSP, 36(11): 1783–1784, 1988.

quadflat

[Byr94] J. S. Byrnes. Quadrature mirror filters, low crest factor arrays, functions achieving optimal uncertainty principle bounds, and complete orthonormal sequences — a unified approach. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 1(??):261–266, 1994. ISSN 1063-5203.

Chui:NM-58-569

[CD91] C. K. Chui and H. Diamond. A general framework for local interpolation. Num. Math, 58:569–581, 1991.

Croisier:ICISS-PCS-76

[CEG76] A. Croisier, D. Esteban, and C. Galand. Perfect channel splitting by use of interpolation/decimation/tree decomposition techniques. Presented at the Int. Conf. Inform. Sci. Syst., Patras, Greece, 1976.

Carpenter:AO-4919

[CG87a] G. A. Carpenter and S. Grossberg.

ART2: Self-organisation of stable category recognition codes for analog input patterns. Appl. Optics, 26:4919–4930, December 1987.

Carpenter:MPA-37-54

[CG87b] G. A. Carpenter and S. Grossberg.

A massively parallel architecture for a self-organising neural pattern recognition machine. Computer Vision, Graphics and Image Processing, 37:54–115, 1987.

Combes:WTMPS-91

[CGT89] J. M. Combes, A. Grossmann, and Ph. Tchamitchian, editors. Wavelets: Time-Frequency Methods and Phase Space. Springer-Verlag, New York, NY, USA, second (1991) edition, 1989.

Chu:ASSP-33-203

[Chu85] Peter L. Chu. Quadrature mirror filter design for an arbitrary number of equal bandwidth channels. IEEE Trans. ASSP, 33(1): 203–218, February 1985.

Chui:MAI-90-11

[Chu90a] C. K. Chui. Construction and applications of interpolation formulas. In Haussmann and Jetter [HJ90], pages 11–23.

Chui:CCS-90-137

[Chu90b] C. K. Chui. Vertex splines and their applications to interpolation of discrete data. In Dahmen et al. [DGM90], pages 137–181.

Chu:LDCCM-92

[Chu92a] K. C. Chu. Lossless data compression circuit and method. U.S. Patent No. 5,150,430, September 1992.

Chui:IW-92

[Chu92b] C. K. Chui, editor. An Introduction to Wavelets. Academic Press, New York, 1992.

Chui:WTTA-92

[Chu92c] C. K. Chui, editor. Wavelets: A Tutorial in Theory and Applications. Academic Press, New York, 1992.

Czulada:MTR-23905wp

[CJ81] J. B. Czulada and J. P. Jacobs. Computerized aids for C^{3} countermeasures decision making. Working Paper 23905, Mitre Corporation, November 1981.

Chiang:AMO-91

[CKS91] T. S. Chiang, G. Kallianpur, and P. Sundar. Propagation of chaos and the McKean-Vlasov equation in duals of nuclear spaces. Appl. Math. and Optimization, 24:55– 83, 1991.

Coifman:CRASP-312-259

[CM91] R. R. Coifman and Y. Meyer. Remarques sur l’analyse de fourier à fenêtre. C.R. Acad. Sci. Paris, 312:259–261, 1991.

Coifman:WTA92-153

[CMW92] R. R. Coifman, Y. Meyer, and M. V. Wickerhauser. Wavelet analysis and signal processing. In Ruskai et al. [RBC^{+}92], pages 153–178.

Cohen:WTA92-105

[Coh92] A. Cohen. Wavelets and digital signal processing. In Ruskai et al. [RBC^{+}92], pages 105–121.

Chui:NMAT-92-a

[CQ92] C. K. Chui and E. Quak. Wavelets on a bounded interval. In Braess and Schumaker [BS92]. To appear.

Chui:TAMU-92-264

[CS92a] C. K. Chui and X. L. Shi. N× oversampling preserves any tight affine-frame for odd N. CAT 264, Texas A&M University, 1992.

Chui:TAMU-92-267

[CS92b] C. K. Chui and X. L. Shi. Bessel sequences and affine frames. CAT 267, Texas A&M University, 1992.

Chui:TAMU-90-230

[CSW92] C. K. Chui, J. Stöckler, and J. D.

Ward. Compactly supported boxspline wavelets. CAT 239, Texas A&M University, 1992.

Chair:AES-22-98

[CV86] Z. Chair and P. K. Varshney.

Optimal data fusion in multiple sensor detection systems. IEEE Trans. Aerospace Electronic Systems, AES-22(1):98–101, January 1986.

Crawford:R-2572-1-AF

[CW85] G. Crawford and C. Williams. The analysis of subjective judgment matrices. Technical Report R-2572-1-AF, Rand Corp., 1985.

Chui:TAMS-330-903

[CW92] C. K. Chui and J. Z. Wang. On compactly supported spline wavelets and a duality principle. Trans. Amer. Math. Soc., 330: 903–915, 1992.

Cybenko:TuftsCS-88

[Cyb88] G. Cybenko. Approximation by superpositions of a sigmoidal function. Technical report, Tufts University, Department of Computer Science, October 1988.

Cybenko:MCSS-2-303

[Cyb89a] G. Cybenko. Approximation by superpositions of a sigmoidal function. Math. Control, Signals and Sys., 2:303–314, 1989.

Cybenko:MCSS-89-2

[Cyb89b] G. Cybenko. Approximation by superpositions of a sigmoidal function. Math. Control, Signals and Sys., 2:303–314, 1989.

DARPA:NNS-90

[DAR90] DARPA. Neural Network Study. AFCEA International Press, Fairfax, VA, 1990.

Daubechies:CPAM-41-909

[Dau88] I. Daubechies. Orthonormal bases of compactly supported wavelets. Comm. Pure Appl. Math., 41:909– 996, October 1988.

Daubechies:TLW-92

[Dau92] I. Daubechies. Ten Lectures on Wavelets. Number 61 in CBMS/NSF Series in Applied Math. SIAM Publ., 1992.

Davis:ZW-64-359

[Dav83] B. Davis. On Brownian slow points. Z. für Wahrscheinlichkeitstheorie, 64:359–367, 1983.

Deming:OC-86

[Dem86] W. E. Deming. Out of the Crisis. MIT, Center for Advanced Engineering, Cambridge, MA, USA, 1986.

ED:TA-92

[Des92] Electronic Design. Technology advances, January 1992.

Daubechies:JMPS-27-1271

[DGM86] I. Daubechies, A. Grossman, and Y. Meyer. Painless nonorthogonal expansions. J. Math and Physics, 27(5):1271–1283, May 1986.

Dahmen:CCS-90

[DGM90] W. Dahmen, M. Gasca, and C. A.

Micchelli, editors. Computation of Curves and Surfaces. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1990.

Dahmen:MP-Freie-Berlin-92

[DK92] W. Dahmen and A. Kunoth. Multilevel preconditioning. Freie Univ. Berlin preprint, 1992.

Davis:AP-13-779

[DP85] B. Davis and E. Perkins. Brownian slow points: the critical case. Annals of Probability, 13:779–803, 1985.

Dahmen:WAMPO-92

[DPS92] W. Dahmen, S. Prössdorf, and R. Schneider. Wavelet approximation methods for pseudodifferential operators. Preprint, 1992.

Drury:CRASP-271-162

[Dru70] S. W. Drury. Sur les ensembles de Sidon. C.R. Acad. Sci. Paris, 271: 162–163, 1970.

Duffin:TAMS-72-341

[DS52] R. Duffin and S. Schaeffer. A class of nonharmonic Fourier series. Trans. Amer. Math. Soc., 72: 341–366, 1952.

Durk:MTR-9140

[DS83] Fay S. Durk and Richard W. Safford. Computer models for assessing the effectiveness of C^{3}CM. Technical Report 9140, Mitre Corporation, November 1983.

Dvoretzky:IJM-1-212

[Dvo63] A. Dvoretzky. On the oscillation of the Brownian motion process. Israël J. Math., 1:212–214, 1963.

Dahlke:Aachen-91-WGM

[DW91] S. Dahlke and I. Weinreich. Wavelet-Galerkin methods: An adapted biorthogonal wavelet basis. Aachen preprint, 1991.

Dahlke:Aachen-92-WBPO

[DW92] S. Dahlke and I. Weinreich. Wavelet bases adapted to pseudodifferential operators. Aachen preprint, 1992.

Dyer:CCW88

[Dye88] Jim Dyer. Command and control warfare, 16 November 1988.

Esteban:CASSP-77-91

[EG77] D. Esteban and C. Galand. Application of quadrature mirror filters to split-band voice coding schemes. In Proc. IEEE Int. Conf. Acoust., Speech, Signal Processing, pages 191–195. Hartford, CT, USA, May 1977.

Erdos:MMJ-4-291

[Erd57] Paul Erdős. Some unsolved problems. Michigan Math. J., 4:291– 300, 1957.

Fan:CRASP-309-295

[H89] Fan Ai-Hua. Sur la convergence de séries trigonométriques lacunaires presque partout par rapport à des produits de Riesz. C.R. Acad. Sci. Paris, 309:295–298, 1989.

Frazier:LPT-91

[FJW91] M. Frazier, B. Jawerth, and G. Weiss. Littlewood-Paley Theory and the Study of Function Spaces. CMBS Regional Converence Series 79. American Mathematical Society, Providence, RI, 1991.

Flandrin:WTMPS-91-69

[Fla91] P. Flandrin. Some aspects of nonstationary signal processing with emphasis on time-frequency and time-scale methods. In Combes et al. [CGT89], pages 68–98.

Fowler:NPCSAOA-83

[Fow83] Commander Naval Systems Sea Command Admiral E. B. Fowler. NAVSEA policy clarification for the specification and assessment of operational availability. Memo, February 14 1983.

Franklin:ACA-30-698

[Fra74] J. Franklin. Ambiguities in the Xray analysis of crystal structures. Acta Crystal A, 30:698–702, 1974.

Freud:MZ-78-252

[Fre62] G. Freud. Uber trigonometrische approximation und fouriersche reihen. Math. Zeitschr., 78:252–262, 1962.

Freud:SM-1-87

[Fre66] G. Freud. On Fourier series with Hadamard gaps. Studia Math., 1: 87–96, 1966.

Frueauf:MTR-8428

[Fru81] Alexander Frueauf. Design of C^{3}CM operational tests. Technical Report 8428, Mitre Corporation, September 1981.

Fleming:IUM-30-925

[FV81] W. H. Fleming and M. Viot. Some measure-valued Markov processes in population genetics theory. Indiana Univ. Math. J., 30:925–935, 1981.

Gabor:JIEEE-93-429

[Gab46] D. Gabor. Theory of communication. J. of IEEE, 93:429–457, November 1946.

Gaposhkin:MUS-33-1

[Gap77] V. F. Gaposhkin. A theorem on the convergence almost everywhere of measurable functions, and its applications to sequences of stochastic integrals. Math. USSR Sbornik, 33:1–17, 1977.

Garling:OSU-77

[Gar77] D. J. H. Garling. Sums of Banach space valued random variables. Mimeographed lecture notes, 1977.

Gitlow:DGQ-87

[GG87] Howard S. Gitlow and Shelly J.

Gitlow. The Deming Guide to Quality and Competitive Position. Prentice-Hall, Englewood Cliffs, N.J., 1987.

Gibbons:NSI-71

[Gib71] J. D. Gibbons. Nonparametric Statistical Inference. McGraw-Hill, New York, 1971.

Glowinski:NMASE-90

[GLRT90] R. Glowinski, W. M. Lawton, M. Ravachol, and E. Tenebaum. Wavelet solution of linear and nonlinear elliptic, parabolic, and hyperbolic problems in one space dimension. In SIAM [SIA90].

Gopinath:Aware-92-MTWGM

[Gop 2] R. Gopinath. Moment theorem for the wavelet-Galerkin method. Technical report, Aware, 1991-2.

Greenberg:JPHN-16-628

[Gre88] S. Greenberg. Acoustic transduction in the auditory periphery. J. of Phonetics, 16:628–666, 1988.

Grossberg:BC-23-121

[Gro76a] S. Grossberg. Adaptive pattern classifications and universal recording, I: Parallel development and coding of neural feature detectors. Biological Cybernetics, 23: 121–134, 1976.

Grossberg:BC-23-187

[Gro76b] S. Grossberg. Adaptive pattern classifications and universal recording, II: Feedback, expectation, olfaction and illusions. Biological Cybernetics, 23:187–202, 1976.

Grenander:TFA-58

[GS58] U. Grenander and G. Szegő. Toeplitz Forms and Their Applications. Berkley University Press, 1958.

Gorman:NN-1-75

[GS88] R. P. Gorman and T. J. Sejnowski. Analysis of hidden units in a layered network trained to classify sonar targets. Neural Networks, 1: 75–89, 1988.

Gonzalez:DIP-87

[GW87] R. Gonzalez and P. Wintz. Digital Image Processing. Addison-Wes-ley, Reading, MA, USA, 2nd edition, 1987.

Hager:TDS-90

[Hag90] G. D. Hager. Task Directed Sensor Fusion and Planning: A Computational Approach. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1990.

Hall:JQO91

[Hal91] C. R. Hall. RADM Jordan’s question on offensive SEW. Memorandum for the Record, 2 December 1991.

Hellman:PP-11-241

[Hel72] R. P. Hellman. Asymmetry in masking between noise and tone.

Perception & Psychophysics, 11: 241–246, 1972.

Higgins:BAMS-12-45

[Hig85] J. R. Higgins. Five short stories about the cardinal series. Bull. AMS, 12:45–89, 1985.

Hoffman:SM-52-159

[HJ74] J. Hoffmann-Jgensen. Sums of independent Banach space valued random variables. Studia Math., 52:159–186, 1974.

Hoffman:LNM-598-1

[HJ77] J. Hoffmann-Jgensen. Probability in Banach spaces. In Ecole d’été de probabilités de Saint-Flour 1976, volume 598 of Lecture Notes in Mathematics, pages 1– 186. Springer-Verlag, 1977.

Haussmann:MAI-90

[HJ90] W. Haussmann and K. Jetter, editors. Multivariate Approximation and Interpolation. ISNM Series Math. Birkhäuser Verlag, Basel, Switzerland, 1990.

Hruscev:CRASP-292-389

[HKK81] S. V. Hruščev, J.-P. Kahane, and Y. Katznelson. Mouvement brownien et séries de Fourier absolument convergentes. C.R. Acad. Sci. Paris, 292:389–391, 1981.

Hardy:NSV-91

[HKRX91] G. Hardy, G. Kallianpur, S. Ramasubramanian, and J. Xiong. Nuclear space valued stochastic differential equations driven by Poisson random measures. Submitted,

Hecht-Nielsen:N-90

[HN90] R. Hecht-Nielsen. Neurocomputing. Addison-Wesley, Reading, MA, USA, 1990.

Hedges:SMM-85

[HO85] V. Hedges and I. Olkin. Statistical Methods for Meta-Analysis. Academic Press, New York, 1985.

Houdre:PTRF-87-167

[Hou90] C. Houdré. Linear Fourier and stochastic analysis. Prob. Th. Rel. Fields, 87:167–188, 1990.

Houdre:AP

[Hou92] C. Houdré. On the spectral SLLN and pointwise ergodic theorem in L^{α}. Annals of Probability, 1992. To appear.

Hovanessian:ISS-88

[Hov88] S. Hovanessian. Introduction to Sensor Systems. Artech House, Norwood, MA, 1988.

Hruscev:TMIS-155-151

[Hru81] S. V. Hruščev. Teorema men’sova ob ispravlenii i Gaussovskie protsessi. Trudy Matem. Inst. Steklov, 155:151–181, 1981.

Hummel:CourantCS-225

[Hum86] R. Hummel. Representations based on zero-crossings in scale space. Technical Report 225, Courant Inst., Dept. Computer Science, 1986.

Hurt:PRZ-89

[Hur89] N. E. Hurt. Phase Retrieval and Zero Crossings. Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1989.

Heil:SR-31-628

[HW89a] C. Heil and D. Walnut. Continuous and discrete wavelet transforms. SIAM Review, 31:628–666, 1989.

Heil:JSR-31-628

[HW89b] C. Heil and D. Walnut. Continuous and discrete wavelet transforms. SIAM Review, 31:628–666, 1989.

Kahane:JAM-14-235

[IIK65] M. Izumi, S. I. Izumi, and J.-P.

Kahane. Théorèmes élémentaires sur les séries de Fourier lacunaires. J. Anal. Math., 14:235–246, 1965.

NeuralWare:npIIug-89

[Inc89] NeuralWare Inc. NeuralWorks Professional II Users Guide, 1989.

Intrator:FR-92

[IRY92] N. Intrator, D. Reisfeld, and Y. Yeshuran. Face recognition using a hybrid supervised/unsupervised neural network, 1992. Preprint.

Jacobs:RAJC

[Jac] J. P. Jacobs. Rear Admiral Jordan’s concerns. Memorandum.

Jackson:IREL-83-21

[Jac83] Peter S. Jackson. On the simportance of elements and prime plicants of non-coherent systems. IEEE Trans. Reliability, R-32:21– 25, April 1983.

Jaffard:LAMM-90-WFRE

[Jaf90] S. Jaffard. Wavelet methods for fast resolution of elliptic problems. Technical report, LAMM, 1990.

Jerri:PIEEE-65-1565

[Jer77] A. J. Jerri. The Shannon sampling theorem—its various extensions and applications: A tutorial review. Proc. IEEE, 65:1565–1596, 1977.

Juran:JQC-88

[Jur88] J. Juran. Juran’s Quality Control Handbook. McGraw-Hill, New York, 4th edition, 1988.

Kahane:AIF-7-293

[Kah57] J.-P. Kahane. Sur les fonctions moyenne-périodiques bornées. Ann. Inst. Fourier (Grenoble), 7:293– 314, 1957.

Kahane:CRASP-248-2943

[Kah59] J.-P. Kahane. Sur un théorème de Paul Malliavin. C.R. Acad. Sci. Paris, 248:2943–2944, 1959.

Kahane:CRASP-259-2577

[Kah64] J.-P. Kahane. Sur les sommes vectorielles ∑ ±u_{n}. C.R. Acad. Sci. Paris, 259:2577–2580, 1964.

Kahane:CRASP-263-613

[Kah66] J.-P. Kahane. Images browniennes des ensembles parfaits. C.R. Acad. Sci. Paris, 263:613–615, 1966.

Kahane:JAT-2-233

[Kah69] J.-P. Kahane. A metric condition for a circular set to be a set of uniqueness. J. Approx. Th., 2:233– 246, 1969.

Kahane:CRASP-278-331

[Kah74] J.-P. Kahane. Sur l’irrégularité locale du mouvement Brownien. C.R. Acad. Sci. Paris, 278:331– 333, 1974.

Kahane:CRASP-282-431

[Kah76] J.-P. Kahane. Sur les zéros et les instants de ralentissement du mouvement brownien. C.R. Acad. Sci. Paris, 282:431–433, 1976.

Kahane:BLMS-12-321

[Kah80] J.-P. Kahane. Sur les polynômes à coefficients unimodulaires. Bull. London Math. Soc., 12:321–342, 1980.

Kahane:TMHA-II-956

[Kah82] J.-P. Kahane. Quatre leçons sur les homéomorphismes du cercle et les séries de Fourier. In Topics in modern harmonic analysis, volume II, pages 956–990, May–June 1982. Proceedings of a seminar held in Torino and Milano.

Kahane:HAZ-83-67

[Kah83] J.-P. Kahane. Slow points of Gaussian processes. In Conference in harmonic analysis in the honour of Antoni Zygmund, volume I, pages 67–83. Wadsworth, 1983.

Kahane:SRS-85

[Kah85] J.-P. Kahane. Some random series of functions. Cambridge University Press, 2nd edition, 1985.

Kahane:JAT-46-1-51

[Kah86] J.-P. Kahane. Geza Freud and lacunary Fourier series. J. Approx. Th., 46(1):51–57, 1986.

Kahane:CQM-51-175

[Kah87] J.-P. Kahane. Distribution des valeurs des fonctions analytiques Gaussiennes. Colloq. Math., 51: 175–187, 1987.

Kahane:LMSLNS-137

[Kah89] J.-P. Kahane. Random multiplication, random coverings, multiplicative chaos. In Analysis at Urbana I (Urbana 1986-87), volume 137 of London Math. Soc. Lect. Note Series, pages 196–255. Cambridge University Press, 1989.

Kahane:CQM-40-387

[Kah90] J.-P. Kahane. Recouvrements aléatoires et théorie du potentiel. Colloq. Math., 40–41:387–411, 1990.

Kahane:RPR-91

[Kah91] J.-P. Kahane. From Riesz products to random sets. Sendai conference, 1990, 1991.

Kallianpur:COSMEX89-227

[Kal90] G. Kallianpur. Infinite dimensional stochastic differential equations with applications. In Proceedings of the 1989 COSMEX Meeting, pages 227–238. World Scientific, 1990.

Katznelson:PAMS-53-396

[Kat75] Y. Katznelson. On a theorem of Menchoff. Proc. Amer. Math. Soc., 53:396–398, 1975.

Kaveh:ASSP-27

[Kav79] M. Kaveh. High resolution spectral estimation for noisy signals. IEEE Trans. ASSP, 27, 1979.

Kovaszny:IRE-560

[KJ55] L. Kovasznay and H. Joseph. Image processing. In Proc. IRE, volume 43, pages 560–570, 1955.

Kahane:CRASP-292-271

[KK81] J.-P. Kahane and Y. Katznelson.

Homéomorphismes du cercle et séries de Fourier absolument convergentes. C.R. Acad. Sci. Paris, 292:271–273, 1981.

Kahane:SFF-83

[KK83] J.-P. Kahane and Y. Katznelson.

Séries de Fourier des fonctions bornées. In Studies in Pure Mathematics to the memory of Paul Turan, pages 395–410. Birkhäuser Verlag, 1983. Preprint Orsay, 1978.

Kallianpur:SSR-29-285

[KMW90] G. Kallianpur, I. Mitoma, and R. L. Wolpert. Diffusion equations in duals of nuclear spaces. Stochastics and Stochastic Reports, 29: 285–329, 1990.

Kolmogoroff:FM-5-96

[Kol24] A. Kolmogoroff. Une contribution à l’étude de la convergence des séries de Fourier. Fund. Math., 5: 96–97, 1924.

Korner:BLMS-12-219

[Kör80] T. W. Körner. On a polynomial of J.S. Byrnes. Bull. London Math. Soc., 12:219–224, 1980.

Kotelenez:LLN-82

[Kot82] P. Kotelenez. Law of large numbers and central limit theorem for chemical reactions with diffusions. Technical report, U. Bremen, 1982.

Kallianpur:AMO-17-237

[KPA88] G. Kallianpur and V. Perez-Abreu. Stochastic evolution equations driven by nuclear spacevalued martingales. Appl. Math. and Optimization, 17:237–272, 1988.

Krylov:SOVM-16-1233

[KR81] N. V. Krylov and B. L. Rozovskii. Stochastic evolution equations. J. of Soviet Mathematics, 16:1233– 1277, 1981.

Kilmer:AIF-38-63

[KS88a] S. J. Kilmer and S. Saeki. On Riesz product measures; mutual absolute continuity and singularity. Ann. Inst. Fourier (Grenoble), 38(2):63–93, 1988.

Konno:PRF-79-201

[KS88b] N. Konno and T. Shiga. Stochastic partial differential equations for some measure-valued diffusions. Prob. Th. Rel. Fields, 79:201–225, 1988.

Kallianpur:AMO-12-125

[KW84] G. Kallianpur and R. Wolpert. Infinite dimensional stochastic differential equation models for spatially distributed neurons. Appl. Math. and Optimization, 12:125– 172, 1984.

Kallianpur:LNB-70-116

[KW87] G. Kallianpur and R. Wolpert. Weak convergence of stochastic neuronal models. In M. Kimura, G. Kallianpur, and T. Hida, editors, Lecture Notes in Biomathematics, volume 70, pages 116–145. Springer-Verlag, 1987.

Kwapien:LNM-256-157

[Kwa76] S. Kwapien. A theorem on the Rademacher series with vector coefficients. In Probability in Banach spaces, volume 526 of Lecture Notes in Mathematics, pages 157– 158. Springer-Verlag, 1976.

Liu:ITNN-3-477

[LAG92] W. Liu, A. Andreou, and M. Goldstein. Voiced-speech representation by an analog silicon model of auditory periphery. IEEE Trans. on Neural Networks, 3:477–487, 1992.

Levinson:GDT-40

[Lev40] N. Levinson. Gap and Density Theorems. American Mathematical Society, New York, NY, USA, 1940.

Levinson:JMP-25

[Lev47] N. Levinson. The Wiener RMS error criterion in filter design and prediction. J. Math and Physics, 25, 1947.

Levy:PSM-48

[Lév48] P. Lévy. Processus stochastiques et mouvement Brownien. Gauthier-Villars, Paris, France, 1948.

Liandrat:WNPDE-ERCOFTAC-90

[Lia90] J. Liandrat. Wavelets and numerical solution of partial differential equations. ERCOFTAC Bulletin, VII, 1990.

Lippmann:ASSP-87-4

[Lip87] R. P. Lippmann. An introduction to computing with neural nets. IEEE Trans. ASSP, pages 4–22, April 1987.

Littlewood:JLMS-36-307

[Lit61] J. E. Littlewood. On the mean values of certain trigonometric polynomials. J. Lon. Math. Soc., 36: 307–334, 1961.

Lozo:ACNN91-22

[LJN^{+}91] P. Lozo, R. P. Johnson, D. Nandagopal, G. Nussey, and T. Zyweck. An adaptive resonance theory (ART) based neural network approach to classifying targets in infrared images. In Proceedings of the Second Australian Conference on Neural Networks (ACNN’91), pages 22– 25. Sydney, NSW, Australia, 1991.

DeLeeuw:CRASP-285-1001

[LKK77] K. De Leeuw, J.-P. Kahane, and Y. Katznelson. Sur les coefficients de Fourier des fonctions continues. C.R. Acad. Sci. Paris, 285:1001– 1003, 1977.

Lee:MML-4-473

[LM83] Ding Lee and Suzanne T. Mc-Daniel. Wave field computations on the interface: An ocean acoustics model. Math. Modeling, 4: 473–488, 1983.

Lorentz:SWODE-91

[LM91] R. A. Lorentz and W. Madych. Spline wavelets for ordinary differential equations. Preprint, 1991.

Lass:SLH-3-82

[LMNY82] N. Lass, L. McReynolds, J. Northern, and D. Yoder, editors. Speech, Language, and Hearing, volume III: Hearing Disorders. W. B. Saunders Company, Philadelphia, PA, USA, 1982.

Littlewood:AM-49-885

[LO48] J. E. Littlewood and C. Offord.

On the distribution of zeros and a-values of a random integral function. Annals of Mathematics, 49: 885–952, 1948.

Littlewood:I-50-990

[LO49] J. E. Littlewood and C. Offord.

Idem, 50:990–991, 1949.

Liandrat:1CASE-90-83

[LT90] J. Liandrat and Ph. Tchamitchian.

Resolution of the 1D regularized Burgers equation using a spatial wavelet approximation. Technical Report 90-83, 1CASE, December 1990. NASA Contractor Report 18748880.

Ledoux:EDM-23-3

[LT91] M. Ledoux and M. Talagrand.

Probability in Banach spaces. Ergebnisse der Mathematik, 23(3), 1991. 480 pages.

Muchnik:FMS-88

[M^{+}88] I. B. Muchnik et al. Functional Multidimensional Scaling. Nauka, Moscow, USSR, 1988.

Malliavin:CRASP-248-2155

[Mal59] P. Malliavin. Sur l’impossibilité de la synthèse spectrale sur la droite.

C.R. Acad. Sci. Paris, 248:2155– 2157, 1959.

Mallat:ASSP-37-2091

[Mal89a] S. Mallat. Multifrequency channel decompositions of images and wavelet models. IEEE Trans. ASSP, 37(12):2091–2110, December 1989.

Mallat:TAMS-315-69

[Mal89b] S. Mallat. Multiresolution approximation and wavelet orthonormal bases of L^{2}. Trans. Amer. Math. Soc., 315:69–87, September 1989.

Malvar:SPL-92

[Mal92] H. Malvar. Signal Processing with Lapped Orthogonal Transforms. Artech House, Norwood, MA, 1992.

Meyer:OO-90

[Mey90] Y. Meyer. Ondelettes et Opérateurs. Hermann, Paris, 1990.

Marr:PRSL-207-187

[MH80] D. Marr and E. Hildreth. Theory of edge detection. Proc. Roy. Soc. London, B207:187–217, 1980.

Mallat:INFT-38-617

[MH92] S. Mallat and W. L. Huang. Singularity detection and processing with wavelets. IEEE Trans. Inf. Theory, 38(2):617–643, March 1992.

Mitoma:MFSK-35-185

[Mit81] I. Mitoma. Martingales of random distributions. Mem. Fac. Sci. Kyushu Univ., 35:185–197, 1981.

Mitoma:AP-11-989

[Mit84] I. Mitoma. Tightness of probabilities on C([0, 1]; S′) and D([0, 1]; S′). Annals of Probability, 11:989–999, 1984.

Motll:LAE-5-714

[MM79] V. V. Motll and I. B. Muchnik.

Linguistic analysis of experimental curves. Proc. IEEE, pages 714– 736, 1979.

Maurey:SM-58-45

[MP76] B. Maurey and G. Pisier. Séries de variables aléatoires vectorielles indépendantes et géométrie des espaces de Banach. Studia Math., 58:45–90, 1976.

Marcus:LNS50-78

[MP78] M. B. Marcus and G. Pisier. Necessary and sufficient conditions for the uniform convergence of random trigonometric series. In Lecture notes series 50. Aarhus, Denmark, 1978.

Miller:INFT-18-241

[MT72] J. H. Miller and J. B. Thomas.

Detectors for discrete time signals in non-Gaussian noise. IEEE Trans. Inf. Theory, IT-18(3):241– 250, March 1972.

Murai:CRASP-287-931

[Mur78] T. Murai. Une remarque sur la distribution des valeurs des séries de Taylor aléatoires. C.R. Acad. Sci. Paris, 287:931–934, 1978.

Murai:JLMS-24-480

[Mur81] T. Murai. Value distribution of random Taylor series in the unit disc. J. London Math. Soc., 24: 480–494, 1981.

Morishita:KYB-11-154

[MY72] I. Morishita and A. Yajima. Analysis and simulation of networks of mutually inhibiting neurons. Kybernetik, 11:154–165, 1972.

Muchnik:IAH-74

[MZ74] I. B. Muchnik and N. Zavalishkin. Image Analysis and Human Perception. Nauka, Moscow, USSR, 1974.

Mallat:CourantCS-92

[MZ92] S. Mallat and Z. Zhang. Matching pursuits with time-frequency dictionaries. Technical report, Courant Inst., Dept. Computer Science, November 1992.

NL

[Nat] National lampoon. Some issues more than others.

Newman:IJ-8-419

[NB70] D. J. Newman and A. Beck. Linear search problems. Israel J., 8:419– 429, 1970.

Newman:PAMS-16-1287

[New65] D. J. Newman. An l_{1} extremal problem for polynomials. Proc. Amer. Math. Soc., 16:1287–1290, 1965.

Newman:AM-138-241

[NF77] D. J. Newman and L. Flatto. Theorems on random coverings. Acta Math., 138:241–264, 1977.

Newman:DM-4-57

[NK73] D. J. Newman and A. G. Konheim.

A note on growing binary trees. Discrete Math., 4:57–63, 1973.

Offord:SM-41-71

[Off72] C. Offord. The distribution of the values of a random function in the unit disc. Studia Math., 41:71–106, 1972.

USGAO:MPU-91

[Off91] United States General Accounting Office. Management Practice: U.S. Companies Improve Performance Through Quality Efforts. United States Government Printing Office, Washington, DC, USA, 1991.

Olevskii:DAN-238-799

[Ole78] A. M. Olevskii. Sučectvovanie funktsii s neustranimymi osobennostjami Carlemana. Dokladi Akad. Nauk SSSR, 238(4):796– 799, 1978.

Olevskii:DAN-256-284

[Ole81] A. M. Olevskii. Zamena peremennoy i absojutnaja shodimost pjada Fourier. Dokladi Akad. Nauk SSSR, 256:284–287, 1981.

Ostheimer:PFP-93

[Ost93] G. Ostheimer. Parallel Functional Programming for Message Passing Multiprocessors. Ph.d. thesis, University of St. Andrews, 1993.

Orey:PLMS-28-174

[OT74] S. Orey and S. J. Taylor. How often on a Brownian path does the law of the iterated logarithm fail? Proc. London Math. Soc., 28(3): 174–192, 1974.

Papoulis:FIA-62

[Pap62] A. Papoulis. The Fourier Integral and its Applications. McGraw-Hill, New York, 1962.

Papoulis:ASSP-29

[Pap81] A. Papoulis. Maximum entropy and spectral estimation: A review. IEEE Trans. ASSP, 29, 1981.

Papoulis:SR-27

[Pap85] A. Papoulis. Levinson’s algorithm, Wold’s decomposition, and spectral estimation. SIAM Review, 27, 1985.

Papoulis:PRV-91

[Pap91] A. Papoulis. Probability, Random Variables, and Stochastic Processes. McGraw-Hill, New York, 3rd edition, 1991.

Patterson:PR-65-195

[Pat44] A. L. Patterson. Ambiguities in the X-ray analysis of crystal structures. Phys. Rev., 65(5):195–201, 1944.

Perrin:LA-14

[Per14] J. Perrin. Les Atomes. F. Alcan, Paris, France, fourth edition, 1914.

Perkins:ZW-64-369

[Per83] E. Perkins. On the Hausdorff dimension of the Brownian slow points. Z. für Wahrscheinlichkeitstheorie, 64:369–399, 1983.

Perlovsky:NNS-87

[Per87] L. I. Perlovsky. Multiple sensor fusion and neural networks. In DARPA Neural Network Study. MIT/Lincoln Laboratory, Lexington, MA, USA, 1987.

Perrier:WTMPS-91-269

[Per91] V. Perrier. Towards a method for solving partial differential equations using wavelet bases. In Combes et al. [CGT89], pages 269–283.

Peyriere:AUSJM-91

[Pey91] J. Peyrière. Almost everywhere convergence of lacunary trigonometric series with respect to Riesz products. Australian J. Math., 1991.

Pickles:IPH-88

[Pic88] J. Pickles. An Introduction to the Physiology of Hearing. Academic Press, New York, 2nd edition, 1988.

Piranashvili:TPA-12-647

[Pir67] P. A. Piranashvili. The problem of interpolation of random processes. Theor. Probability Appl., 12:647– 657, 1967.

Pisier:CRASP-286-671

[Pis78a] G. Pisier. Ensembles de Sidon et processus Gaussiens. C.R. Acad. Sci. Paris, 286:671–674, 1978.

Pisier:SEB-78

[Pis78b] G. Pisier. Les inégalités de Khintchine-Kahane d’après C. Borell. In Séminaire sur la géométrie des espaces de Banach 1977–78. Ecole Polytechnique, 1978.

Pisier:IJM-34-38

[Pis79] G. Pisier. A remarkable homogeneous Banach algebra. Israël J. Math., 34:38–44, 1979.

Pisier:BAMS-8-87

[Pis83] G. Pisier. Arithmetic characterizations of Sidon sets. Bull. AMS, 8:87–89, 1983.

Pisier:VCB-89

[Pis89] G. Pisier. The volume of convex bodies and Banach space geometry. Cambridge University Press, 1989.

Pfeiffer:JASA-58-867

[PK75] R. R. Pfeiffer and D. O. Kim. Cochlear nerve fiber responses: Distribution along the cochlear partition. J. Acoustic Soc. Amer., 58:867–869, 1975.

Prignot:DCE-87

[Pri87] P. Prignot. Dichotomie du cotype pour les espaces invariants. Technical Report 87-02, I1-I50, Publications Mathématiques d’Orsay, 1987.

Paley:CAMS-34

[PW34a] R. E. A. C. Paley and N. Wiener. Fourier transforms in the complex domain. In Coll. Amer. Math. Soc. American Mathematical Society, 1934.

Paley:RTC-34

[PW34b] R. E. A. C. Paley and N. Wiener. Fourier transforms in the complex domain. American Mathematical Society, Rhode Island, 1934.

Paley:MZ-37-647

[PWZ33] R. E. A. C. Paley, N. Wiener, and A. Zygmund. Notes on random functions. Math. Zeitschr., 37:647–668, 1933.

Paley:PCP-26-337-458

[PZ30] R. E. A. C. Paley and A. Zygmund. On some series of functions. In Proceedings of the Cambridge Philosophical Society, volume 26, pages 337–357, 458–474, 1930.

Paley:CPS-28-190

[PZ32] R. E. A. C. Paley and A. Zygmund. On some series of functions. In Proceedings of the Cambridge Philosophical Society, volume 28, pages 190–205, 1932.

Ruskai:WTA-92

[RBC^{+}92] M.-B. Ruskai, G. Beylkin, R. Coifman, I. Daubechies, S. Mallat, Y. Meyer, and L. Raphael, editors. Wavelets and Their Applications. Jones and Bartlett, Boston, 1992.

Rumelhart:LIR-PDP-88

[RHW88] D. Rumelhart, G. Hinton, and R. Williams. Learning internal representations by error propagation. In Rumelhart and McClelland [RM88].

Rider:DMJ-42-759

[Rid75] D. Rider. Randomly continuous functions and Sidon sets. Duke Math. J., 42:759–764, 1975.

Rumelhart:PDP-88

[RM88] David Rumelhart and G. McClelland, editors. Parallel Distributed Processing: Explorations in the Microstructure of Cognition. MIT PressCambridge, MA, 1988.

Ryll:SM-13-30

[RN53] C. Ryll-Nardzewski. D. Blackwell’s conjecture on power series with random coefficients. Studia Math., 13:30–36, 1953.

Reibman:AES-23-24

[RN87] A. R. Reibman and L. W. Nolte. Optimal decision performance of distributed sensor systems. IEEE Trans. Aerospace Electronic Systems, AES-23(1):24–30, January 1987.

Rosenblatt:PN-59

[Ros59] F. Rosenblatt. Principles of Neurodynamics. Spartan Books, 1959.

Rosenblatt:SADM-3-343

[RS82] J. Rosenblatt and P. D. Seymour. The structure of homometric sets. SIAM J. Alg. Disc. Meth., 3:343– 350, 1982.

Rosenfeld:ICOMP-20-562

[RT71] A. Rosenfeld and M. Thurston. Edge and curve detection for visual scene analysis. IEEE Trans. Computers, C-20:562–569, 1971.

Rudin:JMM-9-203

[Rud60] W. Rudin. Trigonometric series with gaps. J. Math. Mech., 9:203– 227, 1960.

Rioul:WSP-91

[RV91] O. Rioul and M. Vetterli. Wavelets and signal processing. ASSP Magazine, October 1991.

Schroeder:ICASSP-85-937

[SA85] M. Schroeder and B. Atal. Code excited linear prediction (CELP): high quality speech at very low bit rates. Proc. ICASSP, pages 937– 940, 1985.

Simoncelli:PIEEE-78-652

[SA90] E. Simoncelli and E. Adelson. Nonseparable extensions of quadrature mirror filters to multiple dimensions. Proc. IEEE, 78 (4):652–664, April 1990.

Saakian:MUS-79-597

[Saa79] A. A. Saakian. Integral’nie moduli gladkosti i koeffitsienti Fourier u superpozitsii funktsii. Math. USSR Sbornik, pages 597–608, 1979.

Sabins:78

[Sab78] F. F. Sabins. Remote Sensing Principles and Interpretation. W.H. Freeman, San Francisco, New York, 1978.

Saffari:CRAS-304-127

[Saf87] B. Saffari. Structure algèbrique sur les couples de Rudin-Shapiro. problème extrémal de Salem sur les polynômes à coefficients ±1. C.R. Acad. Sci. Paris, 304:127– 130, 1987.

Schroeder:JASA-66-1647

[SAH79] M. R. Schroeder, B. S. Atal, and J. L. Hall. Optimizing digital speech coding by exploiting masking properties of the human ear. J. Acoustic Soc. Amer., 66:1647– 1652, 1979.

Salem:TAMS-56-32

[Sal44] R. Salem. Sets of uniqueness and sets of multiplicity 2. Trans. Amer. Math. Soc., 56:32–49, 1944. See § 13.

Salem:AVM-1-353

[Sal50] R. Salem. On singular monotonic functions whose spectrum has a given Hausdorff dimension. Arkiv für Math., 1:353–365, 1950.

Salem:OM-67

[Sal67] R. Salem. Oeuvres mathématiques. Hermann, Paris, 1967.

Schelkunoff:BST-39-80

[Sch60] S. A. Schelkunoff. A mathematical theory of linear arrays. Bell Sys. Tech. J, 39:80–107, 1960.

Schay:IBMJ-6-246

[Sch62] G. Schay. Approximate methods for a multiqueuing problem. IBM J. of Research and Development, 6:246–250, 1962.

Schay:IBMJ-7-350

[Sch63] G. Schay. On a queueing problem arising in recirculating memories. IBM J. of Research and Development, 7:350–353, 1963.

Schuknecht:PE-74

[Sch74] H. Schuknecht. Pathology of the Ear. Harvard University Press, 1974.

Schroeder:PIEEE-63-1332

[Sch75] M. R. Schroeder. Models of hearing. Proc. IEEE, 63:1332–1350, 1975.

Schroeder:FCPL-91

[Sch91] Manfred R. Schroeder. Fractals, Chaos, Power Laws: Minutes from an Infinite Paradise. W.H. Freeman, San Francisco, New York, 1991.

Schonberger:ODY-20-16

[Sch92] Richard J. Schonberger. Total quality management cuts a broad swath—through manufacturing and beyond. Organizational Dynamic, 20(4):16–28, Spring 1992.

Smith:ASSP-38-1446

[SE90] M. Smith and S. Eddins. Analysis/synthesis techniques for subband coding. IEEE Trans. ASSP, 38(8):1446–1456, August 1990.

Shapiro:EPP-51

[Sha51] H. S. Shapiro. Extremal problems for polynomials and power series. Sc.M. thesis, Massachusetts Institute of Technology, 1951.

Shamma:JASA-78-1612

[Sha85a] S. Shamma. Speech processing in the auditory system I: The representation of speech sounds in the responses of the auditory nerve. J. Acoustic Soc. Amer., 78(5):1612– 1621, November 1985.

Shamma:JASA-78-1622

[Sha85b] S. Shamma. Speech processing in the auditory system II: Lateral inhibition and the central processing of speech evoked activity in the auditory nerve. J. Acoustic Soc. Amer., 78(5):1622–1632, November 1985.

NMASE-90

[SIA90] Proc. 9th International Conference on Num. Methods in Appl. Sciences and Engr. SIAM Publ., Philadelphia, 1990.

Siegman:JASA-78-659

[SKL85] W. L. Siegmann, G. A. Kriegsmann, and D. Lee. A wide-angle three-dimensional parabolic wave equation. J. Acoustic Soc. Amer., 78:659–664, 1985.

Skolnik:RH-70

[Sko70] M. Skolnik. Radar Handbook. McGraw-Hill, 1970.

Specht:ICNN88-525

[Spe88] D. F. Specht. Probabilistic neural networks for classification, mapping, or associative memory. In Proceedings of International Conference on Artificial Neural Networks (ICNN-88), volume 1, pages 525–532, 1988.

SPIE:CSC-91

[SPI91] Boston, MA, USA, 1991.

Simmons:1992:TBE

[SSD92] J. A. Simmons, P. A. Saillant, and S. P. Dear. Through a bat’s ear. IEEE Spectrum, 29(3):46–48, March 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Steinhaus:MZ-21-408

[Ste29] H. Steinhaus. Uber die wahrscheinlichkeit dafür, daßder konvergenzkreis einer potenzreihe ihre natürliche grenze ist. Math. Zeitschr., 21:408–416, 1929.

Shannon:MTC-49

[SW49] C. E. Shannon and W. Weaver.

The Mathematical Theory of Communication. University of Illinois Press, 1949.

Stein:IFA-71

[SW71] E. Stein and G. Weiss. Introduction to Fourier Analysis on Euclidean Spaces. Princeton University Press, Princeton, 1971.

Salem:AM-91-245

[SZ54] R. Salem and A. Zygmund. Some properties of trigonometric series whose terms have random signs. Acta Math, 91:245–301, 1954.

Szidon:MA-97-675

[Szi27] S. Szidon. Verallgemeinerung eines satzes über die absolute konvergenz von Fourierreihen mit Lücken. Math. Ann., 97:675–676, 1927.

Taylor:AIF-24-195

[Tay74] S. J. Taylor. Regularity and irregularities on a Brownian path. Ann. Inst. Fourier (Grenoble), 24:195– 204, 1974.

Thomson:JASA-74-1848

[TC83] D. J. Thomson and N. R. Chapman. A wide-angle split-step algorithm for the parabolic equation.

J. Acoustic Soc. Amer., 74(1848-1854), 1983.

Tchamitchian:WTMPS-91-263

[Tch91] Ph. Tchamitchian. About wavelets and elliptic operators. In Combes et al. [CGT89], pages 263–268.

Tremain:CELP-88-491

[TCW88] T. E. Tremain, J. P. Campbell, and V. C. Welch. A 4.8 kbps code excited linear predictive coder. In Proceedings of the Mobile Satellite Conf., pages 491–496, May 1988.

Tappert:ICA74-csl

[TH74] F. D. Tappert and R. H. Hardin. Computer simulation of longrange ocean acoustic propagation using the parabolic equation method. In Eighth International Congress on Acoustics. London, 1974.

Townshend:81

[Tow81] John R. G. Townshend. Terrain Analysis and Remote Sensing. George Allen & Unwin, London, UK, 1981.

Tsitsiklis:IEEEDC-86-232

[Tsi86] J. N. Tsitsiklis. On threshold rules in decentralized detection. In Proc. IEEE Conf. Decision Control, pages 232–236, December 1986.

Thomopoulos:AES-23-644

[TVB87] S. C. A. Thomopoulos, R. Viswanathan, and D. Bougoulias. Optimal decision fusion in multiple sensor systems. IEEE Trans. Aerospace Electronic Systems, AES-23(5): 644–653, September 1987.

Urick:PUS-83

[Uri83] Robert J. Urick. Principles of Underwater Sound. McGraw-Hill, New York, 3rd edition, 1983.

Vaidyanathan:ASSP-35-476

[Vai87] P. P. Vaidyanathan. Theory and design of M-channel maximally decimated quadrature mirror filters with arbitrary M, having the perfect-reconstruction property. IEEE Trans. ASSP, 35(4): 476–492, April 1987.

Vaidyanathan:MFB-92

[Vai92] P. P. Vaidyanathan. Multirate Filter Banks and Wavelets. Prentice-Hall, Englewood Cliffs, N.J., 1992.

VanTrees:DEM-I-68

[Van68] H. L. VanTrees. Detection, Estimation and Modulation Theory, volume I. John Wiley & Sons, New York, London, Sydney, 1968.

Vaidyanathan:ITE-89-181

[VD89] P. P. Vaidyanathan and Z. Doganata. The role of losslessness in modern digital signal processing. IEEE Trans. Education, pages 181–197, August 1989.

Veldhuis:BRASC-92-86

[Vel92] R. N. J. Veldhuis. Bit rates in audio source coding. IEEE Journal on Selected Areas in Communications, 10(1), 1992.

Vetterli:SP-6-97

[Vet84] M. Vetterli. Multidimensional subband coding: Some theory and algorithms. Signal Processing, 6:97– 112, April 1984.

Vetterli:ITA-35-356

[Vet87] M. Vetterli. Multirate filter banks. IEEE Trans. ASSP, 35:356–372, 1987.

Vetterli:MCTDTR-92-161

[VU92] M. Vetterli and K. M. Uz. Multiresolution coding techniques for digital television: A review. Multidimensional Systems and Signal Processing, 3:161–187, 1992.

Walsh:AAP-13-231

[Wal81] J. B. Walsh. A stochastic model of neural response. Adv. Appl. Prob., 13:231–281, 1981.

Walsh:LNM-1180-266

[Wal86] J. B. Walsh. An introduction to stochastic partial differential equations. In P. L. Hennequin, editor, Lecture Notes in Mathematics, volume 1180, pages 266–437. 1986.

Warren:AP-82

[War82] R. Warren. Auditory Perception. Pergamon Press, New York, NY, USA, 1982.

Wickerhauser:ETIT-32-712

[WC92] M. V. Wickerhauser and R. R. Coifman. Entropy based methods for best basis selection. IEEE Trans. Inf. Theory, 32:712–718, March 1992.

Weiss:Aware-92-WDET

[Wei 2] J. Weiss. Wavelets and the dynamics of entropy transfer in two dimensional hydrodynamics. Technical report, Aware, 1991–2.

Wickwire:MTR-25134wp

[Wic83] Kenneth H. Wickwire. Useful MOEs for offensive C^{3} countermeasures. Working Paper 25134, Mitre Corporation, December 1983.

Wickerhauser:CSC-91-23

[Wic91] M. V. Wickerhauser. Fast approximate factor analysis. In Curves and Surfaces in Computer Vision and Graphics II: SPIE Proceedings Volume 1610 [SPI91], pages 23– 32.

Wickerhauser:WTTA92-679

[Wic92] M. V. Wickerhauser. Acoustic signal compression with wavelet packets. In Chui [Chu92c], pages 679–713.

Wiener:JMP-2-131

[Wie23] N. Wiener. Differential space.

J. Math and Physics, 2:131–174, 1923.

Wiener:AM-55-117

[Wie30] N. Wiener. Generalized harmonic analysis. Acta Math, 55:117–258, 1930.

Woods:ASSP-34-1278

[WO86] J. Woods and S. O’Neil. Subband coding of images. IEEE Trans. ASSP, ASSP-34:1278–1288, 1986.

Young:CEP-74

[YC74] T. Y. Young and T. W. Calvert.

Classification, Estimation and Pattern Recognition. Elsevier Publishing Co., 1974.

Young:INFS-80

[You80] Robert M. Young. An Introduction to Nonharmonic Fourier Series. Academic Press, New York, 1980.

Yserentant:NM-58-163

[Yse90] H. Yserentant. Two preconditioners based on the multilevel splitting of finite element spaces. Num. Math, 58:163–184, 1990.

Yao:ITCT-14-404

[YT67] K. Yao and J. Thomas. On some stability and interpolatory properties of nonuniform sampling expansions. IEEE Trans. Circuit Theory, 14:404–408, 1967.

Yang:INFT-38-824

[YWS92] X. Yang, K. Wang, and S. A. Shamma. Auditory representations of acoustic signals. IEEE Trans. Inf. Theory, 38(2):824–839, March 1992.

Zahedi:IF-86-96

[Zah86a] F. Zahedi. The analytic hierarchy process—A survey of the method and its applications. Interfaces, 16 (4):96–108, 1986.

Zahedi:SEP-86-347

[Zah86b] F. Zahedi. A simulation study of estimation methods in the analytic hierarchy process. Socio-Economic Planning, 20(6):347–354, 1986.

Zahedi:COR-87-395

[Zah87] F. Zahedi. Qualitative programming for selection decisions. Computers and Operations Research, 14(5):395–407, 1987.

Zaknich:UWA91-cem

[ZdA91] A. Zaknich, C. deSilva, and Y. Attikiouzel. Comparative evaluation of the multi-class PNN and ART2 neural networks for classification of infra-red ship images— Parts I and II. In The University of Western Australia Research Report to the Guided Weapons Division. University of Western Australia, 1991.

Zucker:PR7

[ZH84] S. W. Zucker and R. A. Hummel. Receptive fields and the representation of visual information. In Proc. 7th Int. Conf. Pattern Recognition, July 1984.

Zweig:JASA-89-1229

[Zwe91] G. Zweig. Finding the impedance in the organ of corti. J. Acoustic Soc. Amer., 89:1229–1254, 1991.

Zygmund:FUM-16-90

[Zyg30] A. Zygmund. On the convergence of lacunary trigonometric series. Fund. Math., 16:90–107, 1930.

Zygmund:TAMS-34-435

[Zyg32] A. Zygmund. On lacunary trigonometric series. Trans. Amer. Math. Soc., 34:435–446, 1932.

Zygmund:ALSS-6-80

[Zyg33] A. Zygmund. On the continuability of power series. Acta Litt. Sci. Szeged, 6:80–84, 1933.

Zygmund:TS-I-II-68

[Zyg68] A. Zygmund. Trigonometric Series I and II. Cambridge University Press, Cambridge, U.K., second edition, 1968.

Zygmund:SPO-89

[Zyg89] Antoni Zygmund. Selected papers of Antoni Zygmund. Kluwer Academic Publishers Group, 1989. 3 volumes.